Dmitri needs 5 liters of a 43% solution of sulfuric acid for a research project in molecular biology. He has two supplies of sulfuric acid solution: one is an unlimited supplyo f the 40% solution and the other an unlimited supply of the 45% solution. How many liters of each solution should Dmitri use?
Found 2 solutions by CharlesG2, ewatrrr:
Answer by CharlesG2(834)

Dmitri needs 5 liters of a 43% solution of sulfuric acid for a research project in molecular biology. He has two supplies of sulfuric acid solution: one is an unlimited supply of the 40% solution and the other an unlimited supply of the 45% solution. How many liters of each solution should Dmitri use?


% A * amount A + %B * amount B = %C * amount C
40% * A + 45% * B = 43% * 5 liters
let B = 5 - A
0.40 * A + 0.45 * (5 - A) = 0.43 * 5
0.4A + 2.25 - 0.45A = 2.15
-0.05A = -0.1
A = 2 liters of the 40%
B = 5 - A = 5 - 2 = 3 liters of the 45%

Answer by ewatrrr(24785)


Hi

Let x represent the 45% solution, then (5-x) would represent the 40% solution
.40(5-x) + .45x = .43*5L
solving for x
.05x = .43*5 - .40*5L
.05x = .03*5 = .15L
x = 3 L of 45% solution and (5-3) or 2L of 40% solution should be used by Dmitri
checking our answer
.40*2 + .45*3 = .08 + 1.35 = 2.15